# Incorporating Magnetic Field Characteristics into EUV-Based Automated Segmentation of Coronal Holes

### v1

@@ 6 Conclusions and Future Work
Through the incorporation of magnetic field information, this new formulation of the ACWE algorithm is able to simultaneously consider region homogeneity in EUV and unipolarity in the magnetic field when delineating between CHs and other solar features. By relying on the fact that CHs appear as dark, homogeneous regions in EUV and as unipolar regions within the magnetic field, this method is able to produce more robust segmentations by retaining CH observations while simultaneously reducing the total area of false positive detections such as surrounding QS (seen in CR 2099) and filaments (seen in CR 2133).
@@
Furthermore, we note that these benefits are retained even with aggressive spatial rescaling of both EUV and HMI magnetogram data to one-eighth of their original size (or $512\times 512$ pixels). This rescaling offers two benefits. First, like with the prior formulation of ACWE for CH detection (Grajeda et al., 2023), the ability to produce meaningful segmentations at a reduced spatial resolution greatly increases algorithm efficiency. Second, we note that operating at a reduced spatial resolution provides an automatic reduction of filament false-positive detections due to the fact that the thin structure of these filaments increases the likelihood that dark pixels will not be retained at the lower resolution, thus eliminating them from the initial threshold-based seed outright. Even when not eliminated, however, the fact that meaningful segmentations are produced at a reduced spatial resolution ensures that filaments that are misidentified in the initial seed are still constrained, thus minimizing overall contamination. With the addition of seed filtering to eliminate non-unipolar regions, filament contamination can be further reduced without compromising the identified CH regions.
@@
With this noted, there are still several areas of potential improvement. First, we note that the seed filtering process herein proposed can eliminate CHs near disk edge that would otherwise be correctly segmented. Addressing this issue may require a variable unipolarity threshold that takes into account the location of the region with respect to disk center. Second, we note that this method relies on line-of-sight magnetic field data, and hope to explore the use of an estimate of radial magnetic field data in future research to determine if this can further improve segmentations near the limb. Finally, we note that the previous formulation of ACWE (Grajeda et al., 2023) provided the ability to quantify the uncertainty of detection based on the homogeneity of any given region compared to adjacent CH regions identified in the initial seeding process. We hope to reintroduce this ability in a future implementation of this expanded ACWE algorithm, leveraging this algorithm’s ability to minimize inclusion of QS regions in order to minimize change of target cases and better characterize the CH region.
@@ Acknowledgements
The authors gratefully acknowledge the support of NASA grant 80NSSC20K0517 which helped support this work.
@@
The data used in this work are publicly available. The GitHub repository at https://github.com/DuckDuckPig/CH-QUACK contains code to download the dataset used in this paper, the base ACWE segmentation code, and code to replicate all experiments described herein.

### v2

@@ 6 Conclusions and Future Work
Through the incorporation of magnetic field information, this new formulation of the ACWE algorithm, QUACK, is able to simultaneously consider region homogeneity in EUV and unipolarity in the magnetic field when delineating between CHs and other solar features. By relying on the fact that CHs appear as dark, homogeneous regions in EUV and as unipolar regions within the magnetic field, QUACK is able to produce more robust segmentations, retaining CH observations while simultaneously reducing the total area of false positive detections. In particular, QUACK reduced the presence of the surrounding QS (seen in CR 2099) and filaments (seen in CR 2133 and in the comparison to ACWE using the Reiss et al. (2023) dataset).
@@
Furthermore, we note that these benefits are retained even with aggressive spatial rescaling of both EUV and HMI magnetogram data to one-eighth of their original size (or $512\times 512$ pixels). This rescaling offers two benefits. First, the ability to produce meaningful segmentations at a reduced spatial resolution greatly reduces computation time. Second, operating at reduced spatial resolution provides a reduction of filament detections since the thin structure of filaments increases the likelihood that dark pixels will not be retained at the lower resolution, eliminating them from the initial threshold-based seed outright. Even when not eliminated, however, the fact that meaningful segmentations are produced at a reduced spatial resolution ensures that filaments misidentified in the initial seed are still constrained, minimizing overall contamination. With the addition of seed filtering to eliminate non-unipolar regions, filament contamination can be further reduced without compromising identified CH regions. This is evident in the results commensurate with Hofmeister et al. (2017); Heinemann et al. (2019) without the application of post-hoc processing to remove filament contamination as needed in those methods.
@@
There are still several areas of potential improvement. First, the seed filtering process herein proposed can eliminate CHs near disk edge that would otherwise be correctly segmented. Addressing this issue may require a variable unipolarity threshold that takes into account the location of the region with respect to disk center. Second, this method relies on LOS magnetic field data, and we hope to explore the use of an estimate of radial magnetic field data to determine if this can further improve segmentations near the limb. Third, the broader relationship between initial seeding parameter and the ideal unipolarity cutoff for seed filtering should be explored to allow for reasonable unipolarity thresholds at values beyond standard seed of $\alpha=0.3$. Finally, we note that the previous formulation of ACWE (Grajeda et al., 2023) provided the ability to quantify the uncertainty of detection based on the homogeneity of any given region compared to adjacent CH regions identified in the initial seeding process. We hope to reintroduce this ability in a future implementation QUACK, leveraging QUACK’s ability to minimize inclusion of QS regions to minimize change of target and better characterize the CH region.
@@ Acknowledgements
The authors gratefully acknowledge the support of NASA grant 80NSSC20K0517, NASA Internal Scientist Funding Model (ISFM) awards to the Wang-Sheeley-Arge (WSA) model team, the Center for Helioanalytics (CfHA), the NSF Division of Atmospheric and Geospace Sciences grant AST-0946422 and multi-messenger astronomy supplemental funding to NSO, and NASA competed Heliophysics Internal Scientist Funding Model (ISFM), all of which helped support this work.
@@
The data used in this work are publicly available. The GitHub repository at https://github.com/DuckDuckPig/CH-ACWE contains code to download the primary dataset used in this paper. The paper Reiss et al. (2023) provides a link to the second dataset. The GitHub repository at https://github.com/DuckDuckPig/CH-QUACK/tree/main/FinalPipeline/DatasetTools contains code to download the final dataset used in this paper. The GitHub repository at https://github.com/DuckDuckPig/CH-QUACK contains code to download the dataset used in this paper, the base ACWE and QUACK segmentation codes, and code to replicate all experiments described herein.